OpenStudy (anonymous):

okay, this is a really dumb question, but i cant figure it out ,so help would be appreciated! (: 1. What is the slope of the line passing through the points (2, 7) and (–1, 3)? (1 point) a. 2/7 b. 3/4 c.4/3 d.1/3

jimthompson5910 (jim_thompson5910):

Use the slope formula m = (y2-y1)/(x2-x1) m = ( 3- 7)/(-1-2) I'll let you finish

jimthompson5910 (jim_thompson5910):

Tell me what you get

OpenStudy (anonymous):

i know how to do that part, but i get for my answers -4/-3... none of which are the options

jimthompson5910 (jim_thompson5910):

-4/-3 reduces to ????

jimthompson5910 (jim_thompson5910):

you're dividing two negatives which do what

OpenStudy (anonymous):

okay so 2 negs = a postive.. i know that, hold on let me grab a calculator

jimthompson5910 (jim_thompson5910):

no need

jimthompson5910 (jim_thompson5910):

-4/-3 = 4/3 and you're done

OpenStudy (anonymous):

oh lol well thank you!!

jimthompson5910 (jim_thompson5910):

the two negatives cancel each other out and go away

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

can you help me with a couple others?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

What is the correct equation of the line shown below?

jimthompson5910 (jim_thompson5910):

first find the slope

OpenStudy (anonymous):

a. y= 3/2x+ 3 b. y= -3/2x-3 c. y= 2/3x + 3 d. y= -2/3x -3 and okay

OpenStudy (anonymous):

okay i got 0/-4... which would be?

jimthompson5910 (jim_thompson5910):

that's not correct

OpenStudy (anonymous):

crap hold on i got this lol

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

0/4?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

okay help lol

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (3-(-3))/(0-(-4)) m = (3+3)/(0+4) m = ??

jimthompson5910 (jim_thompson5910):

you did 3-3 when it should be 3-(-3), which is the same as 3+3

OpenStudy (anonymous):

6/4.. and okay i see what i did wring now, thank you

jimthompson5910 (jim_thompson5910):

reduce that to get ???

OpenStudy (anonymous):

1/5? i think

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

im so rusty on this ists been awhile :/ 1 1/2?

jimthompson5910 (jim_thompson5910):

6 and 4 both have 2 in common so divide both by 2 to get 6/2 = 3 4/2 = 2 so 6/4 reduces to 3/2

jimthompson5910 (jim_thompson5910):

you're close, but 1 1/2 is in a different (unwanted) form

OpenStudy (anonymous):

oh okay okay i get it now and yeah it is lol

jimthompson5910 (jim_thompson5910):

so the slope is 3/2

jimthompson5910 (jim_thompson5910):

the y-intercept is already given on the graph, and that's the point (0,3)

OpenStudy (anonymous):

yep, and okay so its a? and question... what is that one equation for this is it y=mx+b? or something like that

jimthompson5910 (jim_thompson5910):

it is choice A

OpenStudy (anonymous):

m= slope y= y-intercept

jimthompson5910 (jim_thompson5910):

and yes, it's y = mx+b m = slope b = y-intercept

jimthompson5910 (jim_thompson5910):

known as slope-intercept form

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

yes awesome thank you, i am gonna try these last two on my own if i need help i will post it, thank you so much! i just need a brain refresher

jimthompson5910 (jim_thompson5910):

you're very welcome, glad to be of help

OpenStudy (anonymous):

okay ill post this one.. im a little confused

jimthompson5910 (jim_thompson5910):

alright

jimthompson5910 (jim_thompson5910):

go for it

OpenStudy (anonymous):

The x-intercept of a line is –5 and the y-intercept of the line is –2. What is the equation of the line?

jimthompson5910 (jim_thompson5910):

The x-intercept of a line is –5 means that we have the point (-5,0)

jimthompson5910 (jim_thompson5910):

the y-intercept of the line is –2 means we have the point (0,-2)

jimthompson5910 (jim_thompson5910):

so what we're really doing is finding the equation of the line that goes through (-5,0) and (0,-2)

jimthompson5910 (jim_thompson5910):

the steps will be the same 1) find the slope using the two points and the slope formula to get m. 2) find the y-intercept, b, using y = mx+b

jimthompson5910 (jim_thompson5910):

luckily the y-intercept is already given, so step 2 isn't so bad

OpenStudy (anonymous):

okay so i got -2/5

OpenStudy (anonymous):

for the first step

jimthompson5910 (jim_thompson5910):

good, that's your slope

jimthompson5910 (jim_thompson5910):

m = -2/5

jimthompson5910 (jim_thompson5910):

the y-intercept is already given, and it is (0,-2)

OpenStudy (anonymous):

-5 is y?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

y=-2/5x+-2

jimthompson5910 (jim_thompson5910):

or y = -2/5x - 2

jimthompson5910 (jim_thompson5910):

so you got it

OpenStudy (anonymous):

yayy! awesome! okay one more if you could help me through it?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

What is the slope-intercept form of the equation y – 7 = –5/2 (x + 4)? (1 point)

jimthompson5910 (jim_thompson5910):

first distribute

jimthompson5910 (jim_thompson5910):

y – 7 = –5/2 (x + 4) y – 7 = –5/2x - (5/2)(4) y – 7 = –5/2x - 20/2 y – 7 = –5/2x - 10 Then solve for y

OpenStudy (anonymous):

which one first?

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

you distribute the -5/2 through to each term in the parenthesis

OpenStudy (anonymous):

okay, im kinda lost ... how do you solve for why?

jimthompson5910 (jim_thompson5910):

you solve for y because slope intercept form is y = mx+b

jimthompson5910 (jim_thompson5910):

and you can see that y is isolated in slope intercept form